OpenStudy (anonymous):

Name the radius and the equation of the circle with center at(3,-2) , which also passes through the point (0,2).

OpenStudy (anonymous):

Equation of Circle: (x-h)^2 + (y-k)^2 = r^2. Where Center of Circle = (h,k) Thus h = 3, k = -2, Thus equation of circle: (x-3)^2 + (y+2)^2 = r^2. To find radius of circle, it'll just be the distance from the point on the circumference (the point where the circle passes) to the center of the circle. \[Distance Formula = sqrt ((x1-x2)^2 + (y1-y2)^2)\]